OpenStudy (anonymous):

Choose the equation of the line passing through the point (-2, 4) and perpendicular to y = -negative one halfx + 4.

OpenStudy (anonymous):

Choose the equation of the line passing through the point (3, -2) and parallel to y = 3x - 6. y = 3x + 7 y = 3x - 11 y = -3x + 11 y = -3x - 12

OpenStudy (anonymous):

Ignore the first question

OpenStudy (texaschic101):

sorry...I started doing the first question before I noticed your last post. y = 3x - 6 slope here is 3 A parallel line will have the same slope. Now we will use y = mx + b slope(m) = 3 (3,-2)..x = 3 and y = -2 now we sub, and since we already know the slope(m), we will have to find the y intercept (b). y = mx + b -2 = 3(3) + b -2 = 9 + b -2 - 9 = b -11 = b so our parallel equation is : y = 3x - 11

OpenStudy (anonymous):

Ty for the help